# Deliberative XAI: How Explanations Impact Understanding and Decision-Making of AI Novices in Collective and Individual Settings

### v1

@@ Abstract.
XAI research often focuses on settings where people learn about and assess algorithmic systems individually. However, as more public AI systems are deployed, it becomes essential for XAI to facilitate collective understanding and deliberation. We conducted a task-based interview study involving 8 focus groups and 12 individual interviews to explore how explanations can support AI novices in understanding and forming opinions about AI systems. Participants received a collection of explanations organized into four information categories to solve tasks and decide about a system’s deployment. These explanations improved or calibrated participants’ self-reported understanding and decision confidence and facilitated group discussions. Participants valued both technical and contextual information and the self-directed and modular explanation structure. Our contributions include an explanation approach that facilitates both individual and collaborative interaction and explanation design recommendations, including active and controllable exploration, different levels of information detail and breadth, and adaptations to the needs of decision subjects.
@@ 1. Introduction
A growing number of algorithmic decision-making systems (ADM systems) are deployed and used in the public sector to decide about critical issues, such as employment, migration, and criminal justice (Scott et al., 2022; Bansak et al., 2018; Züger and Asghari, 2023; Chouldechova, 2017). These systems can result in discrimination and damages for decision subjects (Brown et al., 2019; Raji et al., 2022), especially when decision-making is opaque and uncontestable (Ananny and Crawford, 2018; Alfrink et al., 2023; de Fine Licht and de Fine Licht, 2020). The field of explainable AI (XAI) aims to reduce these risks by developing methods to make ADM systems more understandable and facilitate their assessment and control (Langer et al., 2021). But while much of XAI research is focused on individuals, research has shown that learning about and understanding complex topics benefits from collaboration in group settings (Smith et al., 2009; Moshman and Geil, 1998; Nokes-Malach et al., 2015; Navajas et al., 2018). Work on public AI further argues that multi-stakeholder deliberation is an essential step in ensuring public participation in deployment decisions of these systems (Kawakami et al., 2024; Naiseh et al., 2024; Brown et al., 2019), but how explanations can fulfill these demands is an open question in XAI.
@@
Crucially, many people affected by decisions of ADM systems have no technical background in developing or analyzing these systems. That is, a majority of decision subjects, domain experts, and other stakeholders of the public are lay people (Lima et al., 2023; DeVito et al., 2018; Shulner-Tal et al., 2022) or AI novices111We use the term in reference to ”visualization novices” (Burns et al., 2023) and synonymously with ”lay people” (Szymanski et al., 2021; Shen et al., 2020; DeVito et al., 2018).. Explanations for AI novices naturally need to fulfill different requirements than for AI practitioners, as they have to be adapted to their background (Ehsan et al., 2024), interests (Langer et al., 2021), and prior knowledge (Szymanski et al., 2021; Cheng et al., 2019; Schmude et al., 2024), which is why personalization (Shulner-Tal et al., 2022; Conati et al., 2021; Naiseh et al., 2020) and interactivity (Cheng et al., 2019; Bertrand et al., 2023; Guesmi et al., 2023) are central concepts in explanation design. Further, while explanation formats (e.g., interactive, visual, textual) are known to impact AI novices’ understanding of algorithmic systems (Cheng et al., 2019; Bove et al., 2022; Szymanski et al., 2021), the precise effect is inconsistent between studies (Bove et al., 2022) due to contextual factors such as participants’ perceptions of the use case domain. Previous qualitative research identified the information needs of AI novices (Schmude et al., 2024), emphasized user-preference for social and conversational formats (Bertrand et al., 2023), and employed participatory methods to design explanations in collaboration with end-users (Weitz et al., 2024; Lee et al., 2019). This work connects to these insights and draws from disciplines such as CSCW (Convertino et al., 2008; Fiesler et al., 2019), cognitive science (Keil, 2006, 2003), and educational psychology (Smith et al., 2009; Nokes-Malach et al., 2015) to examine how explanations can support understanding and decision-making of AI novices in individual and collaborative settings (Naiseh et al., 2024; Weitz et al., 2024).
@@
Developing this design approach to improve understanding and facilitate deliberation meets multiple challenges. Group composition and dynamics place special demands on explanation design (Naiseh et al., 2021), as explanations must cover a diverse set of information needs (Schmude et al., 2024; Bove et al., 2024) and format preferences (Bertrand et al., 2023). In addition, they must support collaborative interaction (Long et al., 2021), such as sharing and simultaneous processing of information. That is, explanations need to provide comprehensive information but remain clear and navigable, and should foster group interaction but still be suitable for individual settings. We address these challenges by proposing a modular explanation design that spans four information categories from which users can select information according to their needs. Additionally, when validating explanation approaches qualitatively, it is crucial to contact the relevant stakeholder groups, which can be especially challenging with decision subjects in minority groups or difficult situations, who arguably have the highest risk of incurring negative consequences from ADM. To address this, we conducted two focus groups with decision subjects to include their perspectives and voices in developing explanations for AI systems in the public sector (Brown et al., 2019).
@@
This paper explores how explanations can be used in individual and collaborative settings to support AI novices’ learning, interaction, and decision-making regarding the deployment of ADM systems. We present the findings of a task-based interview study with 43 AI novices involving 8 focus groups and 12 single interviews. We propose an explanation design that addresses AI novices’ information needs and applies current best practices in XAI (Byrne, 2023). The design consists of a collection of 36 explanations in the form of question-answer pairs organized into the four categories data, system details, usage, and context, which are further subdivided into subtopics and levels of detail (Figure 1). Participants used these explanations to solve the study tasks and decide about deploying an ADM system (Figure 3). The study’s use case is an algorithm for the prediction of job-seekers’ employability, connecting to previous work on ADM systems in employment (Scott et al., 2022; Niklas et al., 2015; Lopez, 2019). We analyze participants’ self-reported understanding, decision confidence, and perceptions of key information, and conduct a thematic analysis of group interactions with explanations. We discuss explanation design suggestions and the benefits and challenges of using explanations in collaborative settings. In the analysis of our findings, we put particular focus on the following research questions:
@@
Explanations: How does a question-driven, modular explanation design support AI novices’ understanding in individual and group settings?
@@
Deliberation: How do the explanations support group interaction with information and deliberation processes?
The contributions of our work include i) the validation of an explanation design building on current XAI principles, including question-driven and modular design, different levels of completeness and soundness, and suitability for both individual and collective settings; ii) an analysis of the kind and amount of information that participants requested most often and perceived as most important; iii) an analysis of how our explanations support group collaboration in solving tasks and group deliberation in forming opinions about ADM systems; and iv) recommendations for the design and use of these explanations and implications for the incorporation of group settings in XAI.
@@ 2. Background and Related Work
In this section, we briefly describe how our work connects to human-centered AI, research on explanation design for AI novices, and literature on understanding, interaction, and decision-making in groups.
@@ 2.1. Human-centered explainable AI
Explainability is often described as a cornerstone of responsible AI systems222While not every ADM system includes AI, we use both terms as we focus on decision-making systems with machine learning components. (Thiebes et al., 2021), as explanations can enable people to understand (Langer et al., 2021) and contest AI decisions (Alfrink et al., 2023). Similar concepts are embedded in the notion of human-centered AI (Capel and Brereton, 2023), which proposes to 1) build systems based on user observation and stakeholder engagement and 2) empower rather than replace people by making these systems controllable and autonomy-preserving (Shneiderman, 2022; Shin, 2023; Xu, 2019). These measures become especially important in high-risk settings (European Commission, 2021), such as employment (Scott et al., 2022; Flügge, 2021), immigration (Bansak et al., 2018), or criminal justice (Chouldechova, 2017), where erroneous and non-transparent algorithmic decisions can cause severe harm to those affected (Raji et al., 2022). To better understand how explanations can contribute to these principles, human-centered explainable AI places emphasis on the sociotechnical dimensions of explaining AI systems (Ehsan et al., 2023b) in that it assumes that transparency alone is not enough to make AI systems explainable (Ananny and Crawford, 2018), but that explanations need to consider the system’s social context (Wenzelburger et al., 2022), its lifecycle (Dhanorkar et al., 2021) and its different stakeholder groups (Ehsan et al., 2023b). In the context of this work, human-centered explainability is realized by testing and validating a design approach that has been developed based on qualitative exploration of AI novices’ information needs (Schmude et al., 2024). Further, our work aims to examine suitable formats for explanations that can accompany the deployment of AI systems in public institutions (de Fine Licht and de Fine Licht, 2020) and inform approaches of practicing responsible AI in high-risk application settings (Ehsan et al., 2023b).
2.2. Explanation design for AI novices
@@
Explanations for non-technical audiences have different requirements than those for AI practitioners. Explanation methods like LIME (Ribeiro et al., 2016), SHAP (Lundberg and Lee, 2017), and surrogate models (Molnar, 2022) are better suited to experts due to their reliance on previous knowledge and technical literacy, while effective methods for AI novices are still being developed. As transparency does not equal understanding (Ananny and Crawford, 2018), simply making all information about a system available is no valid explanation strategy. Instead, explanations must be selected and designed to be suitable to the recipient’s stakeholder role (Langer et al., 2021), prior knowledge (Schmude et al., 2024), beliefs (Miller, 2019), and explanatory stance (Byrne, 2023; Keil, 2006). Various studies examined how such explanations for AI novices might be realized: Cheng et al. (2019) showed that ”white-box” explanations (transparent models) can improve ”objective”333We use the term in reference to Cheng et al. (2019) and Bove et al. (2022), it means to describe factual or testable understanding. understanding but may overwhelm users, reducing perceived understanding. However, Bove et al. (2022) used similar explanations and found that they had the opposite effect on participant understanding, which they attribute to a difference in the studies’ use case domains (student admission vs. finance). Szymanski et al. (2021) examined how expert and lay users rated explanations of different formats and found that while lay users favored visual explanations, they performed better with textual ones. Other studies underscore this variability, indicating that comprehension varies with personal and demographic factors and domain knowledge (Wang and Yin, 2021; Shulner-Tal et al., 2022; Ehsan et al., 2024) and that personalized explanations could be effective in addressing this (Conati et al., 2021; Shulner-Tal et al., 2022). Further aspects to be considered include the explanation’s purpose (Freiesleben and König, 2023) and the recipient’s familiarity with AI (Kramer et al., 2018). Regarding the information needs, previous qualitative research outlined that AI novices value information about the context and intention of a system’s deployment (Schmude et al., 2024; Kawakami et al., 2024) as well about the responsible institution (Brown et al., 2019). Notably, Lee et al. (2019) and Weitz et al. (2024) conducted participatory workshops to design explanations with end-users in the public sector, finding that participants valued the collaborative settings and their improved understanding of algorithmic systems. Drawing from these approaches, this work focuses on the design and qualitative validation of explanations such that they address the information needs of AI novices (Schmude et al., 2024) and support their understanding (Langer et al., 2021) and decision-making (Bertrand et al., 2023) in individual and collective settings.
@@
Our explanation design draws from various approaches in previous work. The information base used for the explanations was compiled from different sources documenting the development and planned deployment of employment prediction algorithms (Scott et al., 2022; Allhutter et al., 2020), producing an extensive collection of ”scavenged” (Wieringa, 2023) material. To structure this collection, we drew from Lim and Dey (2009)’s work on intelligibility types, Liao et al. (2020)’s concept of question-driven explanation design, and Schmude et al. (2024)’s separation of information categories into data, system details, usage, and context. We further apply Kulesza et al. (2013)’s principles on explanation soundness (fidelity, complexity) and completeness (coverage, density) and draw from works that apply these principles in explanation design (Chatti et al., 2022; Guesmi et al., 2023; Kulesza et al., 2015), by introducing a structure of sub-topics and a hierarchy of explanation levels. This combination of question-driven explanations, different levels of information detail and breadth, and user-controlled selection of information aims to introduce modularity and interactivity (Cheng et al., 2019; Schmude et al., 2023; Guesmi et al., 2023) while allowing for adjusting of explanations to users’ needs (Conati et al., 2021; Shulner-Tal et al., 2022).
2.3. Understanding and decision-making in collaborative settings
@@
Improving understanding is a central goal of explanations (Langer et al., 2021) and enables people to pursue their goals, such as assessing or contesting a system’s decisions (Hoffman et al., 2023). However, understanding can be defined in numerous ways (Grimm, 2019; Baumberger et al., 2017; Zagzebski, 2019; Keil, 2006). This work draws from literature in the learning sciences, cognitive sciences, and XAI to define understanding as i) connecting and applying information, whereas ”knowledge” only stores information (Grimm, 2019; Baumberger et al., 2017), ii) being the attempt to grasp the underlying structure of a phenomenon by way of simplification (Zagzebski, 2019), iii) consisting of several ”facets” that include both the analytical and the emotional connection to information (Wiggins and McTighe, 2005), and iv) rarely being exhaustive, but rather a ”working” mental model that is attained by recognizing and filling gaps until the learner deems it sufficient (Keil, 2006). Due to the challenge of defining and measuring understanding (Sato et al., 2019), recent research has proposed an ”abilities-based” approach (Speith et al., 2024) which defines understanding through the skills the learner possesses, connecting to comparable operationalizations by the learning sciences (Wiggins and McTighe, 2005).
While individual understanding has been the subject of many studies in XAI (Chen et al., 2023; Cheng et al., 2019; Wang and Yin, 2021; Schmude et al., 2023), understanding in group settings has been less explored. However, different disciplines have investigated collaborative understanding in detail: In the cognitive sciences, distributed cognition (Keil, 2003) and outsourced understanding (Keil, 2006) are components for the analysis of explanatory understanding, whereas educational psychology produced insight on the effects of group tutoring (Nokes-Malach et al., 2015; Smith et al., 2009; Moshman and Geil, 1998; Navajas et al., 2018) and comparisons to one-on-one tutoring (Bloom, 1984). Further, research in the field of computer-supported collaborative work provides approaches for the analysis of collaborative settings (Fiesler et al., 2019), group composition (Convertino et al., 2008), and group interactions (Sutcliffe, 2005). Recent research on public AI (Züger and Asghari, 2023) has emphasized the importance of including stakeholders by introducing deliberation early in the design process (Kawakami et al., 2024; Yurrita et al., 2022; Zhang et al., 2023). Further, recent work in XAI has begun to consider how explanations for group interactions could be approached, describing that ”many-to-one” interactions (multiple people interacting with an explanation) will likely differ enormously from ”one-to-one” interactions due to ”complexities in group dynamics, cognitive bias amplification, trust issues within the group, and group-centric evaluation” (Naiseh et al., 2024). Drawing from these conceptualizations, this work aims to explore empirically how explanations can support group deliberation by improving understanding and informing decision-making processes.

### v2

# Better Together? The Role of Explanations in Supporting Novices in Individual and Collective Deliberations about AI

@@ Abstract.
Deploying AI systems in public institutions can have far-reaching consequences for many people, making it a matter of public interest. Providing opportunities for stakeholders to come together, understand these systems, and debate their merits and harms is thus essential. Explainable AI often focuses on individuals, but deliberation benefits from group settings, which are underexplored. To address this gap, we present findings from an interview study with 8 focus groups and 12 individuals. Our findings provide insight into how explanations support AI novices in deliberating alone and in groups. Participants used modular explanations with four information categories to solve tasks and decide about an AI system’s deployment. We found that the explanations supported groups in creating shared understanding and in finding arguments for and against the system’s deployment. In comparison, individual participants engaged with explanations in more depth and performed better in the study tasks, but missed an exchange with others. Based on our findings, we provide suggestions on how explanations should be designed to work in group settings and describe their potential use in real-world contexts. With this, our contributions inform XAI research that aims to enable AI novices to understand and deliberate AI systems in the public sector.
@@ 1. Introduction
A growing number of AI systems111We use the term ‘AI system’ to describe algorithmic systems with machine learning components. The nomenclature follows research on explainable AI (Langer et al., 2021) and research on AI in the context of society (Züger and Asghari, 2023; Collins et al., 2024) and regulation (Panigutti et al., 2023). are deployed in the public sector to decide about critical issues, such as employment, migration, and criminal justice (Scott et al., 2022; Bansak et al., 2018; Züger and Asghari, 2023; Chouldechova, 2017). These systems can have consequences for all stakeholders but tend to have the largest impact on their decision subjects (people the system decides over), such as discrimination or misclassification (Brown et al., 2019; Raji et al., 2022). These harms intensify when decision-making is opaque and uncontestable (Ananny and Crawford, 2018; Alfrink et al., 2023; de Fine Licht and de Fine Licht, 2020). For these reasons, public AI systems should be considered as ‘matters of public interest’ (Züger and Asghari, 2023), meaning that they need to be explainable, justifiable, and open for public deliberation (Kawakami et al., 2024; Naiseh et al., 2024; Brown et al., 2019). Explanations can make AI systems more understandable and easier to assess and control (Langer et al., 2021). Much of explainable AI (XAI) research is focused on individuals, but research has shown that group settings can facilitate the understanding of complex topics (Moshman and Geil, 1998; Nokes-Malach et al., 2015; Navajas et al., 2018). Further, group settings encourage the exchange of views and arguments (Stromer-Galley, 2007; Smith et al., 2009), which are vital when engaging in deliberation (collectively finding a solution to a problem (Habermas, 1991)). XAI has not explored in detail how explanations can be combined with group settings to leverage these benefits. In this paper, we aim to address this gap.
@@
As arguably many people affected by decisions of AI systems have no technical background in developing or analyzing these systems, this work focuses on ‘lay people’ (Lima et al., 2023; DeVito et al., 2018; Shulner-Tal et al., 2022) or ‘AI novices’ (Mohseni et al., 2020). Explanations for AI novices naturally have different requirements than explanations for AI practitioners, as they have different expertise (Ehsan et al., 2024), interests (Langer et al., 2021), and prior knowledge (Szymanski et al., 2021; Cheng et al., 2019; Schmude et al., 2024). Explanation formats (e.g., visual, textual, dialogue) are known to impact AI novices’ understanding (Cheng et al., 2019; Bove et al., 2022; Szymanski et al., 2021) but show inconsistent effects (Bove et al., 2022) due to contextual factors such as participants’ perceptions of the use case domain. A possible solution is the adaptation of explanations with personalization (Shulner-Tal et al., 2022; Conati et al., 2021; Naiseh et al., 2020) and interactivity (Cheng et al., 2019; Bertrand et al., 2023; Guesmi et al., 2023). To this end, previous research has analyzed AI novices’ information needs (Schmude et al., 2024) and developed explanations in collaboration with end-users (Weitz et al., 2024; Lee et al., 2019b). However, designs that can adapt to AI novices and that support their deliberation, especially in groups, are still rare (Chiang et al., 2024; Naiseh et al., 2024). But these explanations are essential to provide opportunities where AI novices can learn about and discuss AI systems and to realize the principles of human-centered AI (engaging stakeholders and empowering people (Shneiderman, 2022)).
@@
Designing explanations that support understanding and deliberation for AI novices in both group and individual settings meets multiple challenges. Group composition and dynamics place special demands on explanation design (Naiseh et al., 2021), as explanations need to cater to a diverse set of information and format preferences (Schmude et al., 2024; Bove et al., 2024; Bertrand et al., 2023). They must further support a joint understanding process and collaborative interactions (Long et al., 2021), such as sharing and combining, all while providing comprehensive information and remaining clear and navigable. We address these challenges by proposing a modular explanation design that spans four information categories (data, system details, usage, and context) from which users can select. Another challenge consists in validating explanation approaches qualitatively with the relevant stakeholder groups. Specifically, XAI research does not always include people from marginalized population groups, who are most likely to be affected negatively as decision subjects (Brown et al., 2019). To address this, we conducted two focus groups with decision subjects to include their perspectives and voices on AI systems in the public sector.
@@
To examine the role of explanations in supporting AI novices’ understanding and deliberation we present the findings of a task-based interview study with 43 participants, involving 8 focus groups and 12 single interviews. For this study, we used an explanation design comprising 36 single explanations in question-answer pairs. These explanations are organized into the four categories data, system details, usage, and context and further assigned to subtopics and levels of detail (Figure 1). Participants used these explanations to solve the study tasks and decide about deploying a public AI system (Figure 3). We used an employment scoring algorithm that connects to previous work on AI systems in employment (Scott et al., 2022; Niklas et al., 2015; Lopez, 2019). Our analysis examines participants’ self-reported understanding, decision confidence, and perceptions of key information. We further conducted a thematic analysis of how participants interacted with explanations in both settings. The following research questions guide the analysis:
@@
Explanations: How does a question-driven, modular explanation design support AI novices’ understanding in groups and individual settings?
@@
Deliberation: How do AI novices use explanations to form opinions and make decisions about AI systems?
Our contributions include i) an explanation design that builds on a question-driven and modular design to accommodate different levels of completeness and soundness and that is suitable for both individual and group settings; ii) an in-depth description of how explanations support participants’ understanding and deliberation processes that identifies salient differences between the two settings; iii) an analysis of which type of explanations participants requested most often and perceived as most important; and iv) recommendations regarding the design and use of explanations in group settings. We envision that this work can provide valuable starting points for future XAI research that aims to connect explanations to deliberation on public AI systems.
@@ 2. Background and Related Work
This section describes how our work is embedded in human-centered explainable AI and outlines the main challenges and approaches to designing explanations for AI novices. It further introduces the two main lenses of analysis to answer our research questions: understanding and deliberation.
@@ 2.1. Human-centered explainable AI
Explainability is often described as a cornerstone of responsible AI systems (Thiebes et al., 2021), as explanations can enable stakeholders such as users and decision subjects to understand (Langer et al., 2021) and contest AI decisions (Alfrink et al., 2023). A similar focus is set by the domain of human-centered AI (Capel and Brereton, 2023), which proposes to build AI systems that 1) are based on user-experience design and stakeholder engagement, and 2) empower rather than replace people by being controllable and autonomy-preserving (Shneiderman, 2022; Shin, 2023; Xu, 2019). These principles become especially important in high-risk settings (European Commission, 2024), such as employment (Scott et al., 2022; Flügge, 2021), immigration (Bansak et al., 2018), or criminal justice (Chouldechova, 2017), where erroneous or non-transparent algorithmic decisions can cause severe harm to those affected (Raji et al., 2022). In response to these risks, the domain of human-centered explainable AI (HCXAI) examines how explanations can contribute to “equitable and ethical Human-AI interaction” (Ehsan et al., 2023b). It assumes that transparency alone is not enough to make AI systems explainable (Ananny and Crawford, 2018), but that explanations need to consider the system’s social context (Wenzelburger et al., 2022), its lifecycle (Dhanorkar et al., 2021) and its different stakeholder groups (Ehsan et al., 2023b). In the context of this work, human-centered explainability is realized by testing and validating a design approach intended to support AI novices in understanding AI systems and deciding about their deployment in public institutions (Züger and Asghari, 2023).
2.2. Designing explanations for AI novices
The majority of people who interact and are involved with AI systems in public institutions are lay people or AI novices, here defined as “users who [might] use AI products in daily life but have no (or very little) expertise on machine learning systems” (Mohseni et al., 2020). Established explanation methods, like LIME (Ribeiro et al., 2016), SHAP (Lundberg and Lee, 2017), and surrogate models (Molnar, 2022) are tailored to experts and require technical knowledge; hence, they do not address the needs of AI novices. To better cover these needs and match information to them, it is necessary to understand how non-experts conceive of AI systems. Previous HCI research has analyzed lay understandings to explore user perception and understanding of several algorithmic systems (DeVito et al., 2018; Eslami et al., 2016). Similarly, XAI research has begun to explore the information needs of AI novices to design suitable explanations for a broader audience (Schmude et al., 2024). However, few explanation designs have been proposed that truly assume the perspectives of AI novices (Szymanski et al., 2021; Cheng et al., 2019). In the following, we summarize current approaches with respect to AI novices’ information needs and current practices of explanation design.
@@ 2.2. Designing explanations for AI novices
Regarding information needs, previous qualitative research outlined that AI novices value information about the context and intention of a system’s deployment (Schmude et al., 2024; Kawakami et al., 2024) as well as about the responsible institution (Brown et al., 2019). In contrast to traditional XAI approaches, which focus on descriptive information about the system’s workings and outputs, explanations for AI novices thus also require normative information, such as justifications (Biran and McKeown, 2017) for design choices. Regarding information coverage, previous work posits that transparency does not equal understanding (Ananny and Crawford, 2018) and that simply making all information about a system available is no valid explanation strategy. Empirical evaluations of this claim showed that “white-box” explanations (transparent models) can improve “objective”222We use the term in reference to Cheng et al. (2019) and Bove et al. (2022), it means to describe factual or testable understanding. understanding but may overwhelm non-expert users and reduce perceived understanding (Cheng et al., 2019). However, later work (Bove et al., 2024) used similar explanations and found that they had the opposite effect on understanding, attributed to a difference in the studies’ use case domains (student admission vs. finance). This indicates that the amount of information should be adaptable to the given context. Regarding explanation format, Szymanski et al. (2021) examined how expert and lay users rated explanations of different formats and found that while lay users favored visual explanations, they performed better with textual ones. Other studies confirm this discrepancy and posit that comprehension varies with demographic factors and domain knowledge (Wang and Yin, 2021; Shulner-Tal et al., 2022; Ehsan et al., 2024). These issues are assumed to be addressed with personalization of explanations (Conati et al., 2021; Shulner-Tal et al., 2022), meaning that they are selected and designed according to the user’s stakeholder role (Langer et al., 2021), prior knowledge (Schmude et al., 2024), beliefs (Miller, 2019), and explanatory stance (Byrne, 2023; Keil, 2006). Further aspects to be considered include the explanation’s purpose (Freiesleben and König, 2023) and the user’s familiarity with AI (Kramer et al., 2018). These approaches guided the conceptual development of the explanation design presented in this study.
@@
We compiled information from different sources documenting employment prediction algorithms (Scott et al., 2022; Allhutter et al., 2020), producing an extensive collection of “scavenged” (Wieringa, 2023) material. To structure this collection, we drew from work on intelligibility types (Lim and Dey, 2009), question-driven explanation design (Liao et al., 2020), and the separation of information categories into data, system details, usage, and context (Schmude et al., 2024). We further applied the principles of explanation soundness (fidelity, complexity) and completeness (coverage, density) (Chatti et al., 2022; Guesmi et al., 2023; Kulesza et al., 2015) by introducing a structure of sub-topics and a hierarchy of explanation levels. This combination of question-driven explanations, levels of detail, and user-controlled selection of information aims to support modularity and interactivity (Cheng et al., 2019; Schmude et al., 2023; Guesmi et al., 2023) as well as the adjustment of explanations to users’ needs (Conati et al., 2021; Shulner-Tal et al., 2022). Section 3.2 describes how these principles were realized in the explanation design.
2.3. Analytical lenses: Understanding and deliberation
In the following, we introduce understanding and deliberation to serve as the main analytical lenses for this paper. Section 3.5 then operationalizes them for the evaluation of the explanation design and settings.
2.3.1. Individual and collaborative understanding of AI systems
@@
Improving understanding of an AI system is the primary purpose of explanations, as understanding is thought to enable assessment (e.g., of a system’s fairness) (Langer et al., 2021) and action (e.g., contestation) (Henin and Le Métayer, 2022) for the system’s stakeholders. However, understanding can be defined in numerous ways (Grimm, 2019; Baumberger et al., 2017; Zagzebski, 2019; Keil, 2006). This work draws from research in learning sciences, cognitive sciences, and explainable AI to define understanding as i) connecting and applying information (Grimm, 2019; Baumberger et al., 2017), ii) being the attempt to grasp the underlying structure of a phenomenon by way of simplification (Zagzebski, 2019), iii) consisting of several “facets” that include both the analytical and the emotional connection to information (explain, interpret, apply, take perspective, empathize, self-reflect) (Wiggins and McTighe, 2005), and iv) being a “working” mental model that is attained by recognizing and filling gaps until the learner deems it sufficient (Keil, 2006). Due to the challenge of defining and measuring understanding (Sato et al., 2019), recent research has proposed an “abilities-based” approach (Speith et al., 2024), connecting to comparable operationalizations by the learning sciences (Wiggins and McTighe, 2005). We examine understanding by analyzing which facets of understanding participants use to answer the study tasks and make a confident deployment decision (Section 3.4).
While individual understanding has been the subject of many studies in XAI (Chen et al., 2023; Cheng et al., 2019; Wang and Yin, 2021; Schmude et al., 2023), understanding in group settings has been less explored (Chiang et al., 2023). We thus draw from disciplines that have investigated collaborative understanding: The cognitive sciences have examined distributed cognition (sharing cognitive load) and outsourcing (Keil, 2006) (delegating understanding) as fruitful mechanisms for collaborative settings, such as the navigation of a ship (Keil, 2003). The prerequisite is that groups achieve “cognitive symbioses with mutually supporting roles” (Keil, 2006), i.e., a constructive working dynamic. Similarly, educational psychology has found that peer discussion (Smith et al., 2009), collaborative reasoning (Moshman and Geil, 1998), and aggregated knowledge (Navajas et al., 2018) leads groups to perform better than individuals on the same tasks. However, whether groups perform well depends on their interactions, which can be described with cognitive and social mechanisms of collaborative success and failure (Nokes-Malach et al., 2015). When groups perform worse than individuals, the associated mechanisms include increased memory load and retrieval disruption (losing train of thought). In contrast, members tend to have established common ground and shared task-related information when they perform better. Thus, while it is not clear from the outset if groups are better for learning than one-on-one settings (Bloom, 1984), their advantages, such as sharing of cognitive load and exchange of views, likely support finding solutions to complex problems and present a valuable testing ground to deliberate deployment of AI systems.
The field of computer-supported cooperative work (CSCW) has long examined collaborative settings about group composition and interactions (Fiesler et al., 2019; Convertino et al., 2008; Sutcliffe, 2005). However, work in XAI has only begun to consider how explanations for group interactions could be approached, describing that “many-to-one” interactions (multiple people interacting with an explanation) will likely differ from “one-to-one” interactions due to “complexities in group dynamics, cognitive bias amplification, trust issues within the group, and group-centric evaluation” (Naiseh et al., 2024). Lastly, previous work in XAI has examined individual versus group understanding in AI-assisted decision-making but surprisingly found little effect on understanding (Chiang et al., 2023). Following up on these findings, we use a ‘triangulation of methods’ (Carter et al., 2014), as described in Section 3.5), to empirically explore and compare the effects of explanations on the understanding processes of AI novices in groups and individual settings.
2.3.2. Deliberating on AI systems
Deliberation, in the sense of informed reasoning and decision-making, is based on understanding (de Fine Licht and de Fine Licht, 2020) and is key in enabling citizens to debate public sector AI proposals and their potential consequences (Kawakami et al., 2024; Züger and Asghari, 2023). Habermas (1991) describes deliberation as the exchange of rational-critical arguments on a problem to the end of finding a solution. These rational-critical arguments are grounded in truth or a shared understanding of reality, are open for judgment, and can be defended. This connection between shared understanding and deliberation is central to our examination of explanations’ effects. Deliberation takes place in many areas that shape politics and life in society (Lupia, 2023). Examples include public referendums that let inhabitants vote on jurisdictional changes (such as Swiss federal and state laws (Swiss Confederation, 2025)), citizen forums addressing matters of public importance (such as water supply in California (Innes and Booher, 2003)), and community-based grassroots formats where citizens support each other (such as the right to repair movement (Collins et al., 2024)). These settings have in common that they involve “social entities made up of people who are in one way or another engaged with their environment” and who use deliberation and productive conflict to negotiate and change policy issues (Hajer and Wagenaar, 2003). While these participation formats are not perfect and potentially incur cognitive biases such as groupthink333Prioritizing group harmony over real argumentation and discussion. (Naiseh et al., 2024; Janis, 1971; Baron, 2005), they create spaces where the general public can gather, discuss, form opinions, and decide on public interests. We argue that AI systems in public institutions constitute such public interests, encapsulated in the term public AI. Züger and Asghari (2023) employ the term to make explicit that AI systems in public institutions must fulfill obligations to prove their benefit. These obligations include being justifiable, equal, open to validation, technically secure, and the result of a deliberation or co-design process. Identifying formats supporting this deliberation on public AI systems is an open research challenge. Prior work has investigated how ’mini-publics’ (Fung, 2003) can be used to support the co-design of algorithmic policy (Lee et al., 2019b) and procedural justice in algorithmic resource allocation (Lee et al., 2019a). HCI research has further shown that participatory formats can connect communities and institutions in public service transformation (Crivellaro et al., 2019). And in XAI, studies showed that group discussions can facilitate the critical analysis of an AI system’s recommendations and that supplying information on both pros and cons of an AI’s recommendation lead to more frequent and more productive group deliberation (Chiang et al., 2023).
However, settings that allow participants to deliberate in person on the deployment of high-stakes public AI are underexplored. We aim to address this gap by implementing mini-publics as focus groups with three different compositions (domain experts, decision subjects, and members of the general public444Participants who were neither directly affected as job-seekers nor were potential users of the system.), thus including stakeholders of different backgrounds and degrees of involvement. We further compare group deliberation processes with those in single interviews, which can be described as “internal deliberation” (Mercier and Landemore, 2012). On this basis, we aim to provide insight into suitable explanation designs and social formats to support deliberation on public AI systems. Section 3.5 describes the concrete analysis approach to this end.